# Benchmarking Computer Science Survey Generation

### v1

@@ Abstract
Scientific survey articles play a vital role in summarizing research progress, yet their manual creation is becoming increasingly infeasible due to the rapid growth of academic literature. While large language models (LLMs) offer promising capabilities for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols. To address this gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for evaluating scientific survey generation in the computer science domain. SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers that serves as the retrieval pool. In addition, we propose an automated evaluation framework that measures generated surveys across four dimensions: information coverage, referencing accuracy, structural organization, and content quality. Our evaluation of diverse LLM-based approaches shows that survey generation remains highly challenging, even for advanced self-reflection frameworks. These findings highlight the complexity of the task and the necessity for continued research [^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities in scientific text generation wang2024autosurvey ; gao2024evaluating ; liang2025surveyx , supporting various stages of academic writing from outlining to full-content generation. Among these applications, automated scientific survey generation stands out as a particularly valuable task. Scientific survey articles (also known as survey papers or literature reviews) synthesize knowledge from a broad collection of academic papers into a structured overview of a research area keele2007guidelines ; pautasso2013ten ; boote2005scholars . Traditionally, writing such surveys is highly labor-intensive: an expert in this field must gather dozens or even hundreds of relevant papers, understand their contributions, identify important themes and trends, and finally craft a well-structured survey paper. This labor-intensive process has become increasingly challenging due to the explosive growth of academic literature in recent years. For example, arXiv now receives over a thousand new computer science papers each day, and the annual volume of submissions has more than doubled from 2019 to 2024 liang2025surveyx . Keeping pace with this scale of information is infeasible for individual researchers, prompting growing interest in automatic survey writing.
Despite the success of LLMs in various generation tasks, generating structured and comprehensive scientific surveys has received relatively little attention. A key reason is the lack of standardized benchmarks and evaluation protocols tailored to the unique demands of this task. Unlike general-purpose generation tasks, scientific survey writing remains difficult to evaluate in a quantitative and standardized way, due to the lack of benchmarks that capture its unique requirements: synthesizing multiple sources, ensuring citation accuracy, and maintaining structural coherence. As a result, current systems are often assessed through manual expert reviews nlpcc2024 , which are costly, time-consuming, and hinder reproducibility and systematic progress.
@@
To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark designed to facilitate reproducible research in scientific survey generation. We formalize the task as a two-stage process: given a topic description in the computer science domain (e.g., Machine Learning for Information Retrieval), the system should first collect a set of relevant papers from a large academic corpus and then synthesize a well-structured survey from them. To support this formulation, SurGE provides the necessary resources and a standardized evaluation framework for both the retrieval and generation stages. Each test instance in SurGE comprises a research topic and its corresponding ground-truth survey, which serves as a high-quality reference for evaluation. These ground-truth surveys are carefully selected from peer-reviewed, high-impact survey papers, based on citation statistics and topical relevance. SurGE dataset consists of two main components: (1) a set of test instances, where each instance includes a topic description, its ground-truth survey, and the full list of cited references in that survey; (2) a large-scale academic corpus containing over one million computer science papers, which serves as the retrieval pool for relevant document selection and coverage analysis.
Beyond dataset construction, we also propose an automated evaluation framework that enables reproducible and fine-grained assessment of generated surveys. This framework scores outputs across four key dimensions: information collection, referencing accuracy, structural organization, and content quality. Grounded in established academic writing guidelines webster2002analyzing ; boote2005scholars ; keele2007guidelines ; pautasso2013ten , this framework enables automatic, scalable evaluation without the need for human annotation, making it well-suited for benchmarking and driving systematic progress in this emerging task.
@@
To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark. These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft. The results reveal that even the state-of-the-art systems struggle with the survey generation task, highlighting its difficulty. For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant references. These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research. We hope that SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
@@
    We introduce SurGE, the first comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
@@
    We propose a fully automated evaluation framework that assesses survey quality across four crucial dimensions: information coverage, citation accuracy, structure, and content.
@@ 2 Task Definition
  • Document Collection: A retrieval module or complex retrieval system identifies a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
@@
The output survey $S$ is expected to (i) follow a logical hierarchical organization, (ii) cover key techniques and trends related to the topic, and (iii) provide accurate citations in the appropriate places using a consistent and correct format.
@@ 3 Dataset Construction
3.1 Ground-Truth Survey Collection and Annotation
To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature. Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews. To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact bornmann2008citation ; and (iii) the publish date must be between 2020 and 2024.
Following the initial filtering process, we further refined the SurGE dataset through expert annotation to ensure the inclusion of only high-quality survey documents. This stage aimed to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers. To this end, we recruited a team of four computer science Ph.D. students as annotators. Note that annotating a paper did not require a close reading of the entire document. Annotators could typically complete the task within 8–9 minutes per paper on average. Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) reference quality, which examines the relevance, diversity, and traceability of cited works.
Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable. In cases of disagreement, the paper was discarded to maintain a conservative quality threshold. To ensure fair and motivated participation, annotators were compensated based on working hours at an average rate of 60 CNY per hour, exceeding the local minimum wage in Beijing. Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances. The resulting score of 0.792 indicates substantial agreement and reinforces the reliability of the quality control process. After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers. The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix C.
3.2 Reference Extraction and Processing
[…]
@@ 3.3 Academic Corpus Construction
A crucial component of the SurGE benchmark is a large-scale academic corpus designed to serve as the retrieval pool for the document collection stage. Our corpus is built entirely from scholarly metadata obtained from the arXiv repository. To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license arxiv_tou .
@@ 3.5 Ethical Considerations and Licensing
Our corpus is constructed exclusively from arXiv-provided descriptive metadata (titles, authors, abstracts, identifiers, categories, and license URIs) harvested via the official API. We do not host or redistribute arXiv PDFs or source files. This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication arxiv_tou . This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata arxiv_submission .
@@ 4 Evaluation Framework
Developing a robust evaluation framework for automated scientific survey generation involves identifying critical dimensions that reflect a survey’s overall accuracy, comprehensiveness, and quality. Through discussions with academic experts, peer reviewers, and experienced survey authors, we established four essential evaluation dimensions:Information Collection Quality, Referencing Accuracy, Structural Quality, and Content Quality. Each generated survey is systematically compared with an expert-written Ground Truth (GT) survey, ensuring alignment with high academic standards. In the following subsections, we deeply analyze each evaluation criterion and subsequently define the corresponding quantitative metrics.
4.1 Information Collection Quality
A fundamental requirement for scientific surveys is the comprehensive collection of relevant literature, covering key works that accurately represent the research field. Omissions or insufficient coverage can severely compromise a survey’s credibility, causing readers to question its thoroughness. Therefore, this aspect is quantitatively assessed through the Coverage Score, measuring the extent to which a generated survey includes key literature identified by expert consensus in the GT survey.
Metric Definition.
The coverage score $\mathcal{C}$ is defined as the recall of references included in the generated survey relative to the GT survey:
$\mathcal{C}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
where ${R}_{GT}$ denotes references in the GT survey, and ${R}_{G}$ denotes references in the generated survey. A higher coverage score indicates more comprehensive inclusion of essential literature.
@@ 4.1 Information Collection Quality
While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications. We therefore interpret the Coverage score not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
4.2 Referencing Accuracy
Referencing Accuracy evaluates how well the references cited in a generated survey align with its topic and content. This assessment is crucial for ensuring that the survey is well-supported and that citations are contextually appropriate. We propose a hierarchical evaluation framework that assesses referencing quality from three perspectives:
  • Document-level: Determines if a cited paper is thematically relevant to the survey’s overall topic.
  • Section-level: Evaluates if a reference is placed within a semantically appropriate section of the survey.
  • Sentence-level: Assesses if a reference directly supports the specific claim made in the sentence where it is cited.
4.2.1 Automated Relevance Judgment.
To automate this evaluation, we employ a Natural Language Inference (NLI) model (cross-encoder/nli-deberta-v3-base) to automatically assess relevance. For each reference, we form a premise-hypothesis pair and use the NLI model to predict the relationship between them. The premise is constructed from the reference’s metadata, while the hypothesis is contextualized by its placement in the generated survey.
Specifically, a reference is judged based on the NLI model’s output scores for “Entailment,” “Neutral,” and “Contradiction”:
  • It is deemed relevant if “Entailment” has the highest score.
  • It is deemed partially relevant if “Neutral” is highest and “Entailment” is the second-highest.
  • Otherwise, it is deemed irrelevant.
Since the survey generation process is grounded in our academic corpus, any reference cited by the model that is not found within this corpus is considered a hallucination. Such references are therefore automatically marked as irrelevant across all three hierarchical levels.
4.2.2 Hierarchical Evaluation Metrics.
1. Document-level Accuracy ($R_{d}$). This metric assesses if each reference is relevant to the survey’s main topic. For a cited reference $r$, we first apply a heuristic: if $r$ is also cited in the ground-truth survey, it is automatically marked as relevant. For all other references, we use the NLI model with the following pair:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the paper: ‘{Survey Title}’.”
2. Section-level Accuracy ($R_{s}$). This metric evaluates the contextual appropriateness of a citation within a specific section. We assess each unique (reference, section) pair. If a citation marker in the text points to a non-existent entry in the bibliography, the pair is marked as irrelevant. Otherwise, we query the NLI model:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract is relevant to the section: ‘{Section Title}’.”
3. Sentence-level Accuracy ($R_{t}$). For the most fine-grained evaluation, this metric measures if a reference directly supports the claim in its citing sentence. We assess each unique (reference, sentence) pair using the NLI model:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the sentence: ‘{Sentence Text}’.”
4.2.3 Metric Calculation.
Based on the automated judgments, we define indicator functions for each reference $r$ at a given level: $\mathbb{I}_{\text{rel}}(r)=1$ if $r$ is relevant, and 0 otherwise; $\mathbb{I}_{\text{half}}(r)=1$ if $r$ is partially relevant, and 0 otherwise. The final accuracy score for each level $x\in\{d,s,t\}$ is calculated as:
$R_{x}=\frac{\left|\{r\in\mathcal{R}_{x}\mid r\text{ is relevant}\}\right|+0.5\cdot\left|\{r\in\mathcal{R}_{x}\mid r\text{ is partially relevant}\}\right|}{\left|\mathcal{R}_{x}\right|},$ (2)
where $\mathcal{R}_{d}$ is the set of all unique references in the generated survey, while $\mathcal{R}_{s}$ and $\mathcal{R}_{t}$ are the sets of all unique (reference, section) and (reference, sentence) citation instances, respectively.

### v2

# SurGE: A Benchmark and Evaluation Framework for Scientific Survey Generation

@@ Abstract
The rapid growth of academic literature makes the manual creation of scientific surveys increasingly infeasible. While large language models show promise for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols. To bridge this critical gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for scientific survey generation in computer science. SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers. In addition, we propose an automated evaluation framework that measures the quality of generated surveys across four dimensions: comprehensiveness, citation accuracy, structural organization, and content quality. Our evaluation of diverse LLM-based methods demonstrates a significant performance gap, revealing that even advanced agentic frameworks struggle with the complexities of survey generation and highlighting the need for future research in this area. [^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
@@ 1 Introduction
The volume of scientific literature has been expanding at an unprecedented rate in recent years. For instance, academic archives like arXiv now receive over a thousand new computer science papers daily, more than doubled between 2019 and 2024 (Liang et al., 2025). This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as the manual collection and synthesis of large volumes of relevant papers is both labor-intensive and time-consuming. Faced with this challenge, there is a growing need for automated systems that can generate survey papers effectively. While recent advancements in LLM-based agents (Yao et al., 2023; Wang et al., 2024c) offer significant promise for automating this task, their full potential is severely hindered by the lack of reliable, scalable, and standardized evaluation benchmarks. Currently, the evaluation of automatically generated surveys largely depends on human-expert reviews, which limit the reproducibility and objectivity of assessments (Tian et al., 2024a). Consequently, progress in this field is difficult to quantify, and the relative merits of different methods remain challenging to compare without standardized, multi-faceted evaluation benchmarks.
@@
To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark that establishes a reproducible standard for the survey generation task, which we formalize as a two-stage process: (1) retrieving relevant papers for a given topic from a large corpus, and (2) synthesizing them into a well-structured survey. To support standardized evaluation of both stages, SurGE provides a comprehensive dataset consisting of two key components. The first is a large-scale academic corpus containing over one million computer science papers to be used as the search and retrieval pool. The second is a collection of test instances, where each instance comprises a research topic (e.g., ”Machine Learning for Information Retrieval”) and its corresponding ground-truth survey. Each ground-truth survey is a high-impact, peer-reviewed paper, chosen based on strict citation criteria and subsequently validated for its quality by expert annotators (see Section 3.1), and is provided along with its full list of cited references.
Beyond dataset construction, another key contribution of this work is a fully automated evaluation framework that enables reproducible and multi-faceted assessment of generated surveys. Grounded in established principles for high-quality survey writing (Webster & Watson, 2002; Boote & Beile, 2005; Keele et al., 2007; Pautasso, 2013), our proposed evaluation framework quantifies a survey’s quality across four key dimensions: comprehensiveness, citation accuracy, structural organization, and content quality. This automated approach eliminates the reliance on subjective and time-consuming human evaluation, thus establishing a scalable and standardized methodology to guide and measure progress in survey generation.
@@
To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark. These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft. The results reveal that even the state-of-the-art systems struggle with the survey generation task. For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant citations. These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research in this area. We believe SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
@@
    We introduce SurGE, a comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
@@
    We propose an automated evaluation framework that assesses survey quality across four crucial dimensions: comprehensiveness, citation accuracy, structure, and content.
@@ 2 Task Definition
  • Document Collection: A retrieval module or complex agentic retrieval system collects a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
@@
To enable standardized benchmarking of the survey generation task, SurGE offers a comprehensive framework that formalizes the process. The benchmark includes 205 expert-selected topics, each paired with a topic description $t$ and a ground-truth survey for evaluation. SurGE also provides a large-scale academic corpus $D$ containing over one million papers, which supports the document retrieval stage. The generated surveys are then evaluated through an automated system, which quantitatively assesses their quality across four key dimensions. Details of the dataset construction process are outlined in Section 3, while the evaluation framework is explained in Section 4.
@@ 3 Dataset Construction
3.1 Ground-Truth Survey Collection and Expert Annotation
To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature. Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews. To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann & Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
Following the initial filtering process, we further refine the SurGE dataset through expert annotation. This process aims to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers. To this end, we recruited a team of four computer science Ph.D. students as annotators. Note that annotating a paper did not require a close reading of the entire document. Annotators could typically complete the task within 8–9 minutes per paper on average. Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) citation quality, which examines the relevance, diversity, and traceability of cited works.
Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable. In cases of disagreement, the paper was discarded to maintain a conservative quality threshold. To ensure fair and motivated participation, annotators were compensated based on working hours at an average rate of 60 CNY per hour, exceeding the local minimum wage in Beijing. Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances. The resulting score of 0.792 indicates substantial agreement and reinforces the reliability of the quality control process. After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers. The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix F. For each of the 205 selected ground-truth surveys, we provide its complete reference list, which serves as the gold standard for evaluating the comprehensiveness of a generated survey. The detailed methodology for extracting and processing these references is described in Appendix A.
@@ 3.2 Academic Corpus Construction
A crucial component of the SurGE benchmark is a large-scale academic corpus designed to serve as the retrieval pool for the document collection stage. Our corpus is built entirely from scholarly metadata obtained from the arXiv repository. To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license (arx, 2025b).
@@ 3.4 Ethical Considerations and Licensing
Our corpus is constructed exclusively from arXiv-provided descriptive metadata (titles, authors, abstracts, identifiers, categories, and license URIs) harvested via the official API. We do not host or redistribute arXiv PDFs or source files. This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication (arx, 2025b). This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata (arx, 2025a).
@@ 4 Evaluation Framework
To comprehensively evaluate the quality of automatically generated scientific surveys, we propose a multi-faceted evaluation framework. This framework assesses survey quality across four crucial dimensions: Comprehensiveness, Citation Accuracy, Structural Quality, and Content Quality. Each generated survey is evaluated against an expert-written Ground Truth (GT) survey. The following subsections define the quantitative metrics for each dimension in detail.
4.1 Comprehensiveness
The comprehensiveness of a scientific survey is a critical quality factor, as the omission of key publications can undermine its credibility and value. To quantify this aspect, we evaluate the Recall of a generated survey’s references against the ground-truth reference lists. Formally, let ${R}_{GT}$ be the set of references in an expert-written GT survey and ${R}_{G}$ be the set of references in our generated survey. Recall $\mathcal{R}$ is defined as:
$\mathcal{R}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
@@ 4.1 Comprehensiveness
While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications (detailed in §3.1). We therefore interpret this metric not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
4.2 Citation Accuracy
Citation accuracy is another critical aspect of a high-quality survey. Each citation in a survey should be thematically relevant to the overall topic of the survey, and it must be contextually appropriate in terms of both the section and sentence in which it appears. To evaluate this aspect, we introduce the metric of Citation Accuracy, which evaluates each citation in the survey across three levels. First, at the document level, we assess whether a cited paper is thematically relevant to the overall topic of the survey. Second, at the section level, we evaluate whether a citation is placed in a semantically appropriate section of the survey. Finally, at the sentence level, we verify whether a citation supports the specific claim made in the sentence where it is cited.
To automate this evaluation, we employ a Natural Language Inference (NLI) model (nli-deberta-v3-base [^3] https://huggingface.co/cross-encoder/nli-deberta-v3-base) to assess the relevance of each citation. An NLI model is designed to determine the logical relationship between two text snippets: a premise and a hypothesis. The model then predicts the relationship between these two components, providing probabilities for the following labels: Entailment, Neutral, Contradiction. Due to its ability to capture semantic relationships, NLI has become a standard method for evaluating the factual consistency and relevance of text generated by LLMs, rendering it an ideal tool for our task.
For our specific evaluation, we implement this three-level check by framing it as a series of NLI tasks. For each citation $r$ (with title $T_{r}$ and abstract $A_{r}$) within the generated survey $S$ (with title $T_{S}$), we construct a set of premise-hypothesis pairs. The premise is consistently formulated using the content of the cited paper, providing the evidentiary basis for the claim. The hypothesis is specifically tailored to assert relevance at each of the three levels (document, section, and sentence). This formulation is structured as follows:
NLI Task Formulation Premise (Consistent for all levels): There is a paper. Title: $T_{r}$”. Abstract: $A_{r}$. Hypotheses (Tailored for each level of granularity): Document-level: The paper titled $T_{r}$” with the given abstract is thematically relevant to the survey titled: $T_{S}$”. Section-level: The paper titled $T_{r}$” with the given abstract is relevant to the section: “Section Title”. Sentence-level: The paper titled $T_{r}$” with the given abstract supports the claim made in the sentence: “Sentence Text”.
The score for each citation unit is calculated via a multi-step process. Let $R$ denote the set of all citation instances in the generated survey. For each citation $r\in R$, we compute a score at each of the three levels: document ($y_{d}(r)$), section ($y_{s}(r)$), and sentence ($y_{t}(r)$). The calculation proceeds as follows. First, we resolve two special cases without querying the NLI model. Any citation $r$ not found in our academic corpus is classified as a hallucination and assigned a score of $y_{x}(r)=0$ at all levels $x\in\{d,s,t\}$. Conversely, any citation $r$ that is present in the ground-truth survey’s bibliography is assigned a document-level score of $y_{d}(r)=1$. For all other cases, the relevance score $y_{x}(r)$ is determined by the NLI model’s output probabilities for the labels Entailment, Neutral, Contradiction. The probabilities are mapped to a final score for each unit as follows:
[…]
Finally, we aggregate the individual citation scores to produce three final metrics for the survey: Document-level Accuracy ($R_{d}$), Section-level Accuracy ($R_{s}$), and Sentence-level Accuracy ($R_{t}$). For each level $x\in\{d,s,t\}$, the score $R_{x}$ is calculated as the mean of the individual citation scores $y_{x}(r)$ over all citation instances $\mathcal{R}$ in the survey:
$R_{x}=\frac{1}{|{R}|}\sum_{r\in{R}}y_{x}(r),\quad x\in\{d,s,t\},$ (3)